### v1

# Neutrino flux sensitivity to the next galactic core-collapse supernova in COSINUS

Abstract
While neutrinos are often treated as a background for many dark matter experiments, these particles offer a new avenue for physics: the detection of core-collapse supernovae. Supernovae are extremely energetic, violent and complex events that mark the death of massive stars. During their collapse stars emit a large number of neutrinos in a short burst. These neutrinos carry 99% of the emitted energy which makes their detection fundamental in understanding supernovae. This paper illustrates how COSINUS (Cryogenic Observatory for SIgnatures seen in Next-generation Underground Searches), a sodium iodide (NaI) based dark matter search, will be sensitive to the next galactic core-collapse supernova. The experiment is composed of two separate detectors which will be sensitive to far away and nearby supernovae. The inner core of the experiment will consist of NaI crystals operating as scintillating calorimeters. These crystals will mainly be sensitive to the Coherent Elastic Scattering of Neutrinos (CENS) against Na and I nuclei. The low mass of the cryogenic detectors enables the experiment a sensitivity to close supernovae below 1 kpc without pileup. The crystals will see up to hundreds of CENS events from a supernova happening at 200 pc. They reside at the center of a large cylindrical 230 T water tank, instrumented with 30 photomultiplier tubes. This tank acts simultaneously as a passive and active shield able to detect the Cherenkov radiation induced by impinging charged particles from ambient and cosmogenic radioactivity. A supernova near the Milky Way Center (10 kpc) will be easily detected inducing 60 measurable events, and the water tank will have a 3 sensitivity to supernovae up to 22 kpc, seeing 10 events. This paper shows how, even without dedicated optimization, modern dark matter experiments will also be able to play their part in the multi-messenger effort to detect the next galactic core-collapse supernova.
1 Introduction
Triggered by its own gravitational collapse, the death of a massive star is marked by the emission of a violent shock-wave (core-collapse supernova) and the birth of new compact objects, such as neutron stars or black holes. These supernovae are some of the most intense sources of neutrinos in our universe, as over a timescale of a few seconds, 99 of the stellar gravitational binding energy is converted into neutrinos with energies on the order of a few tens of MeV [1]. SN1987A was the first supernova from which neutrinos were observed, but only about two dozen events where detected from this astrophysical phenomenon [2]. The rate of galactic core-collapse supernovae has been predicted to be 1.63 0.46 per century, but this rate is not uniform throughout the galaxy [3]. One of the stars closest to earth that is likely to undergo a core-collapse supernova is Betelgeuse which is located at 200 pc [4].
Observation of neutrinos from the next galactic core-collapse supernova will provide critical information about the neutrino driven explosion mechanism, the properties of neutrinos in ultra-dense environments, verify current hydro-dynamical modeling of stars, possibly provide evidence of new physics, and more [1, 5, 6]. As such, in the subsequent decades since the last supernova, many dedicated supernova neutrino observatories have been established [7, 8, 9, 10, 11, 12] or are planned [13, 14, 15, 16, 17] with sensitivities varying over the local galactic group (kpc - Mpc).
The current generation of supernova observatories primarily consists of large (a few m3 to km3), monolithic target detectors, and are mostly sensitive to and . Currently, the largest detectors sensitive to supernova neutrinos are Super-Kamiokande [18] and IceCube [19]. Both predominantly expect to detect through inverse beta decay and some through charged current interactions. However, the detection of all other flavors is important for understanding the stellar mechanics of supernovae. In 2017, Coherent Elastic Neutrino Scattering (CENS) was observed for the first time by the COHERENT experiment [20]. This interaction channel is sensitive to all flavors, has a high interaction cross-section (relative to other neutrino channels), and future experiments, such as RES-NOVA [21, 22], will seek to exploit this channel in the search for supernova neutrinos. In addition, direct detection dark matter experiments have also begun evaluating this channel for their sensitivity to galactic core-collapse supernovae [23, 24, 25, 26].
The COSINUS (Cryogenic Observatory for SIgnatures seen in Next-generation Underground Searches) experiment [27, 28] will operate sodium iodide (NaI) crystals as cryogenic scintillating calorimeters, with the primary purpose of cross-checking the DAMA/LIBRA experiment [29, 30, 31, 32]. COSINUS is located at the Laboratori Nazionali del Gran Sasso (LNGS) under 1400 m of rock overburden (3600 m.w.e) and will feature a dual-channel read-out system of both heat and scintillation light. The dual-channel read-out system will allow for event-by-event discrimination between e-/ interactions and nuclear recoils [33, 34]. The experiment is designed to minimize the amount of background nuclear recoils. This is achieved by choosing a site under a large overburden, developing low-radioactive crystals, ensuring that materials close to the detectors are as radiopure as possible, and passive [35] and active shielding [36]. This effort is designed to give a background rate of less than one nuclear recoil event per kg per year. COSINUS can search for CENS of supernova neutrinos in the nuclear recoil spectrum of the NaI crystals, and since the active shielding is a water Cherenkov detector other interaction channels such as inverse beta decay can be explored.
This paper will investigate the neutrino interaction channels offered by the COSINUS experiment and its sensitivity to the next core-collapse galactic supernova. Section 2 will detail the experimental setup as well as the different phases of the experiment. Section 3 will explore the neutrino detection channels available in both the NaI crystals and the water Cherenkov detector, while Section 4 will discuss the galactic sensitivity of these channels to the next supernova. Finally, Section 5 will summarize the results presented herein.
2 Experimental Setup
The COSINUS experimental facility is built to house a dry dilution refrigerator with 48 superconducting quantum interference devices (SQUIDs) for amplification and read-out of the sensors. Inside the dilution refrigerator are copper boxes that contain the NaI detector modules. During the operation of the cryostat, the NaI detector modules will be operated at temperatures on the order of 10 mK. Besides reducing thermal noise, this allows for the use of transition edge sensors (TESs) made from tungsten films, to measure the phonon signal. A COSINUS NaI detector module consists of two parts: the NaI target equipped with a TES and a silicon beaker. Due to NaI’s low melting point and softness, depositing a TES directly onto the absorber cannot be done. Instead, a TES is deposited on a sapphire wafer and is linked to a gold pad placed on the crystal with a thin gold wire. This setup is called the “remoTES” [37]. A TES is also deposited on the silicon beaker to measure the scintillation light of particle interactions. The design of the beaker allows for maximized light absorption. The combination of light and phonon channel allows for event-by-event particle discrimination, as well as setting the absolute energy scale, which would be impossible with just one channel.
For the first phase of the experiment (referred to as COSINUS-1), one box with 8 modules will be installed. Each module will have one (2.1 cm)3 NaI crystal, for a total mass of 272 g of target material. This configuration will be used for the first year of data taking. Then in the second phase (COSINUS-2), two more boxes are planned to be added. This will bring the total to 24 modules with a total target material mass of 816 g. If more exposure is required, 117 g hexagonal crystals could be installed in each of the 24 modules bringing the total up to 2808 g (COSINUS-Max).
To minimize the background, additional shielding has been installed in the facility. This includes 30 cm of copper which is put inside the dilution refrigerator, above the detector boxes. During operation, there is an additional 8 cm of copper shielding surrounding the cryostat inside the drywell. The experimental setup including both copper shields can be seen in Fig. 1.
A water tank is used as a passive and active shield for the experiment. During operation, the dilution fridge is lowered into a drywell so that the detector modules are surrounded by water without the outside of the cryostat getting wet. This drywell is 3.66 m deep and has a radius of 0.33 m. The water tank itself is 7 m tall with a diameter of 7 m and is filled with ultra-pure water. The active veto is accomplished with 30 photomultiplier tubes (PMTs) arranged inside the tank. The PMTs are 8-inch diameter R5912-30 from Hamamatsu [38]. Their arrangement can be seen on the floor and walls of the water tank in Fig. 1. To account for the physical space of the PMTs and to reduce triggers from ambient gamma rays, a 30 cm dead layer was introduced near the walls of the water tank. Tyvek, a highly reflecting material, is hung with 30 cm gap from the wall providing the dead layer and an improving light collection. The type of Tyvek used is 1082D, with two layers of Tyvek sandwiching a polyethylene layer. The reflectivity of this foil has been measured to be at least 95% [39]. The properties needed for the foil and the positioning of the PMTs were investigated using Monte Carlo simulations. The tank was optimized to be as efficient as possible as a muon veto [36].
2.1 Monte Carlo Modelling of the Water Tank
The geometry of the setup has been implemented using a Geant4 [40, 41, 42] based software toolkit called ImpCRESST [43], which was initially developed within the CRESST [44] dark matter search and later also within COSINUS [36]. ImpCRESST utilizes Geant4 v10.2.3 and root v6-08-06 [45]. In the simulation, the geometry is modeled as described in Sec. 2. A 30 cm dead layer is achieved with a reflective foil that was set to have 95% reflectivity. Thirty PMTs are setup inside the tank with 18 being placed on the bottom and 12 along the wall. This is two more PMTs than what was studied in [36] as the technical detail of the veto changed since that study. Along the bottom they are arranged into two concentric circles of 9 PMTs each, with the inner circle having a radius of 1.5 m and the outer having a radius of 3.2 m. Along the wall they are placed in two rows, evenly spaced, of 6, where the PMTs are equally distributed around the tank. The results of the ImpCRESST simulation consist of information such as particle tracking, energy deposition, particle type, interaction mechanism, and timing information. For this particular study, the propagation of the optical photons was also recorded.
3 Supernova Neutrino Interaction Channels in COSINUS
During a core collapse supernova 31053 ergs [46] of neutrinos will be emitted within a short time window, approximately 10 s. This intensity over a small timescale makes these violent, astrophysical events excellent sources of neutrinos for terrestrial detectors. Interactions with these terrestrial detectors can take many forms (CENS, inverse beta decay, charge current interactions etc.) and their sensitivity is highly target dependent.
3.1 Supernova Neutrino Fluence Models
The exact details of how the explosion takes place and how neutrinos are emitted is not fully understood and quite model dependent [1]. Given this uncertainty this paper evaluates two different neutrino fluence models. The first adopts a quasi-thermal spectrum [47, 48] to describe the differential neutrino fluence in energy:
where corresponds to the supernova’s distance from Earth and to the neutrino energy. Given that experiments can not distinguish between and supernova neutrinos, they are treated as having equal properties . To have a comparison with already measured data, parameters corresponding to the best-fit for SN1987A events [46] were considered: MeV, MeV, MeV and . The total energy of the supernova shared equally between all neutrino flavors, is 1053 erg. For the second model, an explosion with a progenitor simulated by [49] was used. Due to these parameters varying in time within the simulation, only the time-averaged fluence was considered. These fluences can be seen in Fig. 2.
3.2 Neutrinos in the COSINUS Experiment
The COSINUS experiment can be broken down into two primary subsystems: the NaI cryogenic calorimeters and the water Cherenkov detector. Inside these subsystems the dominant interaction channels of impinging supernova neutrinos will be uniquely different.
3.2.1 Neutrino Interaction Channels in NaI cryogenic calorimeters
For neutrinos interacting inside the NaI crystals, CENS is the primary detection channel. Neutrinos will scatter off the nuclei causing recoils and depositing a few keV of energy into the target. The light output of nuclear recoils is quenched when compared to that of e-/ events and can easily be distinguished by the COSINUS detector module setup [34]. The target nuclear recoil background for the experiment will be less than one event per kg per year, so neutrino events corresponding to a galactic core collapse supernova are expected to coincide with a background-free environment. Additionally, CENS is flavor independent and will be insensitive to neutrino oscillation. The differential recoil rate is given as
where is the number of target nuclei and is the neutrino flux summed over all flavors defined by Eqn. 3.1. The minimum energy required for a target nucleus to recoil is
where is the recoil energy of the nucleus, and is the target nucleus mass. The standard model differential cross-section [50, 51, 52] for neutrino scattering can be written as
where is the Fermi constant, and are the number of protons and neutrons of the target nucleus, is the weak mixing angle and is the form factor. The Klein-Nystrand [53] form factor was used for this analysis and is defined as
where , j1 is the spherical Bessel function of order one, is the range of the Yukawa potential (0.7 fm [54]) and . is the root-mean square neutron (proton) radius and is taken as 2.92 fm (2.99 fm) for 23Na [55, 56] and 4.84 fm (4.75 fm) for 127I [57, 56].
An additional channel of neutrino detection available in NaI crystals is charged current (CC) interactions with 127I, defined as
and
The neutrino energy threshold for lepton production will be 1 MeV, and the electron and positron produced will be visible in the e-/ band. Given the large mass of the Te/Xe nuclei most of the neutrino energy becomes the lepton’s kinetic energy ( MeV). Detector efficiencies at such energies will be close to 100. Charged current interactions with or neutrinos are not possible as the energy of the supernova neutrinos will be below the required production threshold. Fig. 3 compares the different interaction cross-section of neutrinos at supernova energies in NaI. The CENS cross-sections are calculated directly from Eqn. 3.4111Where knowledge of physical parameters is required (e.g., ), the values are taken from [58] unless otherwise specified. while the charged current cross-section are taken from SNOwGLoBES [59, 60]. It is clear from Fig. 3 that within the NaI crystals the CENS channel in 127I will dominate the neutrino interactions by more than an order of magnitude. There are other channels of neutrino interactions available for the NaI crystals (neutral current (NC) with 23Na or 127I or charged current with 23Na) but either the interaction cross-section will be too small and no interactions would occur.
3.2.2 Neutrino Interaction Channels in the Water Cherenkov Detector
When a charged particle travels through water at a speed faster than the speed of light in that medium Cherenkov radiation is produced [61]. Whether through primary or secondary production neutrino interactions can produce charged particles of sufficient energy to create this radiation. Electrons and positrons must have a total energy above 0.77 MeV for Cherenkov radiation to be created. Table 1 gives a summary of the possible neutrino-water tank interactions with any secondary reactions also being listed.
Process | Interaction | Detectable | Cherenkov | Production |
Threshold | Threshold | |||
Proton Scattering | No | 1.4 GeV | 0 MeV | |
Electron Scattering | Yes | 0.77 MeV | 0 MeV | |
Inverse Beta Decay | Yes | 0.77 MeV | 1.80 MeV | |
MeV | No | 0.77 MeV | - | |
CC Oxygen | OF | Yes | 0.77 MeV | 15.9 MeV |
CC Oxygen | ON | Yes | 0.77 MeV | 10.9 MeV |
NC Oxygen ([62]) | OO∗ | No | - | - |
16OO | Yes | 0.77 MeV | 6.05 MeV |
The COSINUS water tank is sensitive to neutrino processes showcased in Table 1. Several other processes not listed in Table 1 were considered, but the water tank is insensitive to these channels due to threshold or efficiency constraints. Firstly, proton scattering will not be visible in the water Cherenkov detectors due to its high threshold requirement for the production of Cherenkov radiation (1.4 GeV). Secondly, the subsequent 2.2 MeV gamma produced from the neutron capture is also not detectable due to the low efficiency of the COSINUS water tank at those energies, see Sec. 3.2.3. The recoil of the scattered oxygen atoms in neutral current reactions will not create Cherenkov radiation and is invisible to our detector. Finally, there are many available excited states of 16O during the neutrino neutral current scattering interaction. The first excited state is at 6.05 MeV, corresponding to the minimum required neutrino energy for this process. Most of the higher level excited states will decay through the production of an alpha particle [62] and similar to the proton, this particle will be below the threshold energy of Cherenkov production. The exception will be the first five states which will de-excite through the production of -rays. The energy of the -rays will be between 6.05 - 7.11 MeV, so for the analysis in this study an average energy of 6.58 MeV will be assumed.
For most of the detectable processes222Estimation of electron scattering in the water tank from supernova neutrinos cannot be analyzed using Eqn. 3.8, and will behave similarly to CENS as described in Sec. 3.2.1. Appendix A gives the full details for this process. listed in Table 1 the number of detected supernova events can be given by [63]
where is the number of targets for the interaction , is the supernova neutrino flux defined in Sec. 3.1, is the interaction cross-section, is the detection efficiency and is the threshold energy defined in Table 1. For the energy of supernova neutrinos the cross-sections of the interactions of Table 1 are shown in Fig. 4, with the data taken from SNOwGLoBES [59, 60, 64]. It is clear that for the average expected energy of supernova neutrinos (10 - 20 MeV) inverse beta decay will be the primary channel of neutrino interaction. However, detector efficiency can increase or decrease the prominence of certain interaction channels, which is discussed in the following section.
3.2.3 Water Cherenkov Detector Efficiency
Electrons, positrons and -rays are the detectable products of neutrino interactions within the water tank. As described in Sec. 2.1, the COSINUS water tank was modeled in Geant4 to evaluate the detection efficiency of these particles over a typical supernova energy range (1 - 100 MeV). The simulation was performed by selecting the energy and particle type and uniformly distributing it inside the water tank volume. The particle momentum was isotropic, and the amount of light that arrived in the PMTs was recorded. Different trigger conditions (i.e. PMT threshold, PMTs in coincidence, etc.) could then be evaluated to determine if the particle was detected. The efficiency is determined as the number of detected particles over the total number of particles simulated. This process was then repeated over the above mentioned energy range. Fig. 5 shows the energy dependent efficiency for positrons, electrons and -rays. For this analysis a single photo-electron threshold was used, as this will be the operating threshold of the COSINUS PMTs [36], and the quantum efficiency of the PMTs was selected to be 30, corresponding to the specifications of the manufacturer at the Cherenkov wavelength. Fig. 5 also shows how the efficiency varies depending on the number of PMTs that are required to trigger for an event to be classified as detected. Based on concerns of high-background triggers the COSINUS experiment will require 4 - 6 PMTs to trigger during run-time conditions. The efficiency for detecting positrons and electrons was found to be practically identical, while the -ray efficiency was about 10 less. Specifically, the efficiency for detecting the 2.2 MeV -ray from a neutron capture was found to be 0.071, making this channel unviable for this experiment.
3.3 Neutrino Oscillation
Neutral current scattering processes such as CENS or electron scattering are flavor independent and neutrino oscillation effects will be negligible. However, charged current processes like inverse beta decay, are sensitive to only specific flavors of neutrinos and oscillation effects must be taken into account. Incoming and may change to and , respectively, (or vice-versa) during their propagation to Earth. Following the process described by [65, 66, 67], which considers vacuum oscillations, and the Mikheyev-Smirnov-Wolfenstein effect in the stellar envelope but not Earth-matter effects, the and flux reaching earth can be described by
and
where is the initial neutrino flux defined by Eqn. 3.1 and and are the survival probabilities. The survival probabilities will be dependent on the neutrino mass hierarchy and are given as
and
4 Sensitivity to Core-Collapse Supernova Neutrinos
Through the combination of a large, optically instrumented water tank and modular cryogenic NaI calorimeters, the COSINUS experiment will have sensitivity to a multitude of neutrino interaction channels. These include: CENS, inverse beta decay, neutral and charged current interactions, and electron scattering. The following sections explore how the discovery significance of these channels will change as a function of distance to the supernova, and discuss how event pile-up can limit the sensitivity to near-supernovae.
4.1 Detection in NaI Cryogenic Calorimeters
The COSINUS detector modules, described in Sec. 2, will have the ability to distinguish nuclear recoils from other e-/ interactions. With the target neutron background goal of less than one event per kg per year and the short (10 s) length of supernovae, this will allow for zero background events and thus, direct identification of CENS interactions inside the detector module. A trigger efficiency of 100 in the phonon channel of the NaI module is assumed for this analysis [34]. By integrating over the neutrino energy in Eqn. 3.2, Fig. 6 shows the expected CENS recoil energy spectra for a 400 pc supernova. Due to its higher mass, the 127I isotopes will be the dominant target for CENS interactions, while the 23Na isotopes will be capable of recoils at higher energies. The target threshold for the COSINUS detectors is 1 keVnr, and by integrating the recoil energy spectrum above the threshold the total number of expected CENS events can be determined. For a total detector mass of 1 kg and the recoil energy spectrum shown in Fig. 6, this gives 10.6 3.2 expected events. With zero expected background events during the timescale of a supernova the significance ( [68, 69]) can be characterized as 3.2, or colloquially at a level indicating evidence of a supernova.
Fig. 7 shows the expected number of CENS events, along with the statistical error, for an array of supernova distances for both considered models. Three different target mass phases of the COSINUS experiment: COSINUS-1 (272 g), COSINUS-2 (816 g), and COSINUS-Max (2808 g) are shown. The 3 and 5 significance band, denoting when evidence or discovery of a supernova can be claimed, are also overlaid.
For a 1987A-like supernova the different target mass phases will be able to claim evidence for a supernova at 0.300 kpc, 0.60 kpc, and 1.06 kpc, respectively. Signal pile-up inside the detector module will reduce the total number of events visible during a supernova and is taken into account in Fig. 7. The probability of having a signal pile-up over a specified time window with a signal rate is given by
Assuming an average duration for a supernova of 10 s, that CENS interactions are uniformly distributed in the number of modules for each target mass, and that the time window around a remoTES cryogenic event is 100 ms [37] the probability of a signal event arriving within the same window as another event can be determined. In Fig. 7 the pileup events are removed from the total expected events, resulting in the bending of the CENS event distribution seen at very near distances. Ultimately, due to the modular nature of the COSINUS NaI, the effect of pileup is only noticeable at supernova distance less then 0.1 kpc.
The expected number of CC127I events was also investigated using the cross-section shown in Fig. 3 and Eqn. 3.8. Unfortunately, the number of events will not be significant at any relevant supernova distances. For example, for the COSINUS-2 phase, at the Betelgeuse distance of 0.197 kpc, only 1 event is expected to be observed. However, knowledge of the CC127I cross-section is fairly limited with the COHERENT experiment just recently publishing the first measurement of the energy-inclusive cross-section [70]. Therefore, in the event of a supernova, the e-/ band should still be analyzed in search of this interaction channel.
Although the distance to which the NaI crystals are sensitive is small compared to the size of the local galactic group (100 kpc), these modular detectors are excellent for detecting nearby, core-collapse supernovae, such as the candidates discussed in [71]. This is due to their sensitivity to all flavors of neutrinos, lack of backgrounds in the nuclear recoil band, and low-pileup threshold because of their modular nature.
4.2 Detection in Water Cherenkov Detector
The sheer size of the water tank will allow for a supernova sensitivity to much larger distances than the crystals which reside inside. The number of expected events in the water tank can be determined by applying Eqn. 3.8, the cross-sections in Fig. 4, the efficiencies in Fig. 5 and the effect of neutrino oscillation (discussed in Sec. 3.3) to the available interactions in Table 1. The exception is the electron recoil interactions, which are detailed in Appendix A. For a supernova at 10 kpc, Table 2 gives the number of expected events per interaction type, for normal and inverse ordering scenarios and for a PMT triggering requirement of 4 and 6-fold. Oxygen NC interactions are given as upper limits due to the small number of expected events. The error given is purely statistical, as it was found that systematic errors of the efficiency and required measured parameters were negligible. As expected, it is clear from this table that inverse beta decay events will be the dominant contribution to the total number of neutrino events.
Process | Ordering | PMT-fold | 1987A Events | Events |
Electron | Normal | 4 | 2 1 | 2 1 |
Scattering | Normal | 6 | 1 1 | 1 1 |
Inverse | 4 | 2 1 | 2 1 | |
Inverse | 6 | 1 1 | 1 1 | |
Inverse Beta | Normal | 4 | 34 6 | 41 6 |
Decay | Normal | 6 | 22 5 | 27 5 |
Inverse | 4 | 63 8 | 62 8 | |
Inverse | 6 | 43 7 | 42 6 | |
CC | Normal | 4 | 1 1 | 1 1 |
Oxygen | Normal | 6 | 2 1 | 1 1 |
Inverse | 4 | 1 1 | 1 1 | |
Inverse | 6 | 1 1 | 1 1 | |
CC | Normal | 4 | 1 1 | 1 1 |
Oxygen | Normal | 6 | 1 1 | 1 1 |
Inverse | 4 | 3 2 | 2 1 | |
Inverse | 6 | 3 2 | 1 1 | |
NC Oxygen | Normal | 4 | 1 | 1 |
Normal | 6 | 1 | 1 | |
Inverse | 4 | 1 | 1 | |
Inverse | 6 | 1 | 1 |
Fig. 8(d) shows the number of total expected events in the water tank as a function of supernova distance. Both the normal and inverse ordering are considered, as well as 4 and 6-fold PMT triggering conditions. The 3 significance denoting the evidence of a supernova is also overlaid. Similar to the analysis in Sec. 4.1, the pileup events have been removed from the total with an assumed trigger window of 1 s, based on the tank characterization from [36].
The 3 discovery potential is drawn as a horizontal line and the distance of the important galactic reference points are shown. Signal events that would arrive in the tank during a typical PMT pulse length have also been subtracted from the distribution.
Unlike the NaI calorimeters the water tank will have a natural background rate, which has been characterized in [36] and consists of muon events, ambient radiation and thermionic emissions inside the photocathodes of the PMTs. For a 4(6)-fold PMT triggering requirement the background rate was determined to be 1.05 Hz (0.09 Hz).
The maximum distance at which the water tank could claim evidence of a supernova as a function of supernova model and PMT triggering condition can be seen in Tab. 3. In the most optimistic scenario the water tank will be sensitive up to kpc considering a 1987A-like supernova and inverted ordering. The horizon distance increases by up to when requiring 6 PMTs to be triggered as opposed to 4, this clearly shows how in that configuration the reduction of background is favored despite reducing the number of signal events. The number of PMTs that are required to trigger a signal does reduce the total event rate, however the large decrease in background rate allows the water tank to extend the horizon by a few kpc. This allows the water tank to be sensitive to supernovae just at the edge of the Milky Way galaxy.
Model | 1987A | |||
PMT-fold | 4 | 6 | 4 | 6 |
NO horizon (kpc) | ||||
IO horizon (kpc) |
5 Conclusion
The COSINUS dark matter experiment has been evaluated for its sensitivity to the next galactic core-collapse supernova. Through modular, cryogenic, NaI calorimeters and a large water Cherenkov detector the COSINUS experiment can probe supernovae both near and far in the Milky Way. With CENS the NaI crystals will be sensitive to all flavors of neutrinos from supernovae up to 1 kpc. Their modular nature minimizes the effects of pileup and the ability for event-by-event particle discrimination creates a background free environment during a supernova explosion. Supernova neutrino interactions in the water tank will be dominated by inverse beta decay events and for the assumptions considered in this study the COSINUS water tank could be sensitive to a supernova between 15 and 22 kpc away. Future work will involve integrating the COSINUS experiment into the Supernova Early Warning System [72] and evaluating the directional sensitivity of the water Cherenkov detector. Moreover, as discussed in [47], COSINUS can be able to identify the time of black-hole formation in case of Failed CCSNe with an uncertainty of a few ms, providing a useful alert for the nearby gravitational wave interferometer Virgo to look for the GW counterpart. Finally, the COSINUS experiment, with one large homogeneous detector containing modular cryogenic calorimeters in the center, is distinctly sensitive to almost all of the possible neutrino interactions in matter (CENS, CC, NC, and electron scattering). Although COSINUS was not optimized for the detection of supernova neutrinos, the variety of interaction channels it is sensitive to and the possibility of coincidence analysis could represent a design philosopy to follow for future dedicated supernova neutrino experiments.
Appendix A Electron-Neutrino Scattering
As discussed in Sec. 3.2.2 neutrinos can scatter off the electrons present in the water tank. Traditionally, this interaction can provide some directionality with regard to the location of the supernova. This process is sensitive to all flavors of neutrinos and the differential recoil rate is described as
where is the number of target electrons and is the neutrino flux defined by Eqn. 3.1. Based of conservation of momentum, the minimum neutrino energy to generate a electron recoil of energy can be written as
where is the mass of the electron. From [73] the differential cross-section can be written as
with the coefficients and being different for each flavor of neutrino, see Table 4.
sin | sin | - sin | sin | |
sin | sin | sin | - sin |
For a supernova at 10 kpc the differential recoil spectrum is shown in Fig. 9. Both the normal and inverse ordering scenario are shown, as well the efficiency from Fig. 5 is folded in. By integrating the spectra and multiplying by the water tank mass the total number of electron scattering events can be calculated. Table 5 shows the number of events for a supernova at 10 kpc. Repeating the above process for different supernovae distances a profile of the total number of electron scattering events can be built, shown in Fig. 10. As shown, electron scattering will contribute only a small percentage of the total number of neutrino interactions and will only be relevent for very close supernovae.
Flux | Ordering | PMT Threshold | Total | ||||
1987A | NO | 4 | 1.64 | 0.17 | 0.31 | 0.34 | 2.46 |
1987A | IO | 4 | 1.30 | 0.30 | 0.36 | 0.17 | 2.13 |
1987A | NO | 6 | 0.92 | 0.075 | 0.16 | 0.18 | 1.34 |
1987A | IO | 6 | 0.70 | 0.14 | 0.19 | 0.08 | 1.11 |
NO | 4 | 1.45 | 0.20 | 0.32 | 0.35 | 2.32 | |
IO | 4 | 1.24 | 0.30 | 0.35 | 0.29 | 2.18 | |
NO | 6 | 0.75 | 0.09 | 0.16 | 0.18 | 1.18 | |
IO | 6 | 0.63 | 0.14 | 0.18 | 0.15 | 1.10 |